20. Solve the system of equations by the addition method.
-36(x-6)=35y
6x-6y=-35
Simplify your answer. Type an ordered pair.

Answer by unlockmath(1688)

Hello, First let's rewrite this as:
-36x-35y=-216
6x-6y=-35
Now we can multiply the second equation by 6 to get:
36x-36y=-210
And add this to the first to get:
-71y=-426
Divide -71 into both sides to get:
y=6
Plug this in to find:
x=1/6
Make sense?
RJ
www.math-unlock.com